Can anyone please help me here
Lets assume that the combines population of Anchorage and Tampa was 25,000 in 1995. By 1960 the population of Anchorage had increased by 10% while Tampa had increased by 9%. If the total population increased by 2380 poeple, then what was the population of each city in 1995.

Answer by jainenderkapoor(61)

THE COMBINED POPULATION OF BOTH THE CITIES IS GIVEN FOR YEAR 1995.
LET US ASSUME THAT THE POPULATION OF CITY ANCHORAGE IS x
THEN THE POPULATION OF CITY TAMPA BECOMES 25000-x
POPULATION INCREASE IN ANCHORAGE BY 1960

10% of x
POPULATION INCREASE IN TAMPA BY 1960

9% of (25000 - x)
TOTAL POPULATION INCREASE IS GIVEN BY
10x/100 + 9*(25000-x)/100 = 2380
(10x + 9*25000 - 9x)/100 = 2380
10x - 9x + 225000 = 2380*100
x + 225000 = 238000
x = 238000 - 225000
x = 13000
SO population of ANCHORAGE IN 1995 WAS 13000
AND population of TAMPA IN 1995 WAS 25000 - 13000 = 12000
HOPE THIS IS CLEAR TO YOU. IN CASE YOU NEED MORE HELP YOU CAN CONTACT ME.
I AM AVAILABLE FOR ONLINE TUTORING. my email is kapoorjai1@rediffmail.com